# Particle production, absorption, scattering, and geodesics in a Schwarzschild–Hernquist black hole

### v1

@@ III.1.1 Thermal radiative emission
The contribution to the emitted radiation associated with frequencies around $\omega$ is obtained by restricting attention to the spectral flux within the differential window $\omega,\omega+\mathrm{d}\omega] [$15]. Carrying out this evaluation yields
@@ V Partial wave analysis
Taking into account Eq. (83), the effective potential for the scalar field exhibits a localized barrier that vanishes asymptotically at the event horizon and at spatial infinity. The asymptotic form of the radial wavefunction $\psi(r^{*})$ is therefore constrained: it must be purely ingoing at the horizon, while at infinity it consists of a superposition of incident and reflected waves. Imposing these physical boundary conditions yields  [64, 5]
@@ VI.1 Low–Frequency Regime
At sufficiently low frequencies, the absorption cross section of a black hole is known to converge to the area of its event horizon. This universal behavior has been extensively established in the literature and represents a robust feature of black hole scattering processes [18, 34, 36]. In the long–wavelength regime, where the incident wave probes length scales comparable to or larger than the black hole size, the detailed structure of the spacetime becomes subdominant. As a result, the absorption process is governed primarily by the geometric properties of the horizon, causing the cross section to asymptotically approach the horizon area.
In particular, for scalar field perturbations propagating on a Schwarzschild background, the low–frequency absorption cross section is analytically shown to coincide with the geometrical area of the event horizon. Remarkably, this limit persists across a wide range of scenarios, including rotating black holes, modified black hole parameters, and nontrivial surrounding geometries [56, 11, 65], provided the incident frequency remains sufficiently small.
Consequently, in the low–frequency regime, the absorption cross section can be expressed as
@@
The low–frequency behavior of the absorption cross section for different values of the Hernquist parameters is illustrated in Fig. 7. As expected, when both $r_{\text{s}}$} and $\rho_{\text{s}}$} vanish, the spacetime reduces to the Schwarzschild black hole, for which the low–frequency absorption cross section approaches $\sigma_{\text{low}}\simeq 16\pi M^{2}$. For a fixed value of ${r_{\text{s}}}/M=0.1$, increasing the dimensionless density parameter ${\rho_{\text{s}}}M^{2}$ from 0 to $0.6$ produces only a negligible modification in the low–frequency absorption cross section. This weak dependence is clearly visible in the left panel of Fig. 6, where $\sigma_{\text{abs}}$ remains close to the Schwarzschild case, $16\pi M^{2}$, at the low–frequency regime for different values of ${\rho_{\text{s}}}M^{2}$.
In contrast, the absorption cross section exhibits a more pronounced sensitivity to variations in the scale radius $r_{\text{s}}$}. As shown in the right panel of Fig. 7, increasing ${r_{\text{s}}}/M$ leads to a noticeable enhancement of the absorption cross section. This trend is consistent with the behavior displayed in Fig. 6, where larger values of ${r_{\text{s}}}/M$ correspond to higher absorption cross sections at low frequencies. $\checkmark$
@@ VI.2 High–Frequency Regime
In this section, we investigate the high–frequency limit of the absorption cross section, commonly referred to as the geometric cross section, $\sigma_{\text{geo}}$. In this regime, wave effects become negligible and the absorption process is well described by null geodesic motion. The geometric cross section is defined as
@@
Substituting Eq. (99) into Eq. (98), the critical impact parameter can be obtained, which in turn allows us to compute the high–frequency absorption cross section. The final expression for $\sigma_{\text{geo}}$ reads
@@
Figure 6 displays the total absorption cross section obtained from partial waves with $l=0$ to $l=6$ for different values of the Hernquist parameters. The corresponding geometric cross section, computed from Eq. (101), is shown by dashed horizontal lines. As expected, the total absorption cross section approaches a constant value in the high–frequency regime, confirming its convergence toward the geometric cross section $\sigma_{\text{geo}}$.
[…]
@@ VII Scattering cross section
The observable angular distribution of the scattered radiation is characterized by the differential scattering cross section, which is related to the scattering amplitude by [83]
@@
The total scattering cross section is obtained by integrating the differential cross section over the full solid angle. For a monochromatic wave of frequency $\omega$, this leads to
@@
This expression highlights that the total scattering cross section arises from a coherent superposition of all partial waves, with each contribution weighted by the corresponding phase shift.
VII.1 Partial scattering cross section
@@
To investigate the influence of the Hernquist black hole parameters on the scattering properties, we analyze the angular dependence of the differential scattering cross section for selected multipole numbers. In particular, Figs. 89 display the differential cross section as a function of the scattering angle $\theta$ for several representative values of $\ell$, illustrating how individual partial waves contribute to the overall scattering pattern.
@@
Figure 8: The normalized partial differential scattering cross sections of the scalar wave for $M\omega=1$, $\ell=1,~2$, ${r_{\text{s}}}/M=0.3$ and various different Hernquist parameter ${\rho_{\text{s}}}M^{2}$.
@@
Figure 8 presents the partial differential scattering cross sections for scalar waves with $M\omega=1$ and ${r_{\text{s}}}/M=0.3$, for several choices of ${\rho_{\text{s}}}M^{2}$, comparing the cases $\ell=1$ and $\ell=2$, in the left and right panle, respectively. In both multipole numbers, a clear and consistent pattern emerges: increasing ${\rho_{\text{s}}}M^{2}$ leads to a higher overall amplitude of the partial scattering cross section. While the $\ell=1$ and $\ell=2$ curves display their usual differences in angular sharpness, their dependence on ${\rho_{\text{s}}}M^{2}$ remains uniform, indicating that the effect of this parameter is broadly insensitive to the multipole structure.
@@
Figure 9: The normalized partial differential scattering cross sections of the scalar wave for $M\omega=1$, $\ell=1,~2$, ${\rho_{\text{s}}}M^{2}=0.1$ and various different Hernquist parameter ${r_{\text{s}}}/M$.
@@
Figure 9 displays the partial differential scattering cross sections for scalar waves with fixed parameters $M\omega=1$ and ${\rho_{\text{s}}}M^{2}=0.1$, considering the multipole modes $\ell=1$ and $\ell=2$ in the left and right panels, respectively. For both modes, the qualitative behavior is the same: the partial scattering cross section increases as the Hernquist parameter ${r_{\text{s}}}/M$ becomes larger. Although the $\ell=1$ and $\ell=2$ curves differ in the shape of their angular patterns, they exhibit the same monotonic dependence on ${r_{\text{s}}}/M$.
@@
Figure 10: The normalized partial scattering cross sections for different multipole numbers when ${\rho_{\text{s}}}M^{2}={r_{\text{s}}}/M=0.1$.
@@
Figure 10 displays the scalar wave differential scattering cross sections in the Hernquist black hole spacetime for several multipole numbers, evaluated at ${\rho_{\text{s}}}M^{2}={r_{\text{s}}}/M=0.1$. As the angular momentum mode $\ell$ increases, from $\ell=1$ to $4$, the angular structure of the scattering becomes progressively sharper and more oscillatory, with the dominant contribution concentrated near the forward direction where the cross section reaches its maximum. This behavior is consistent with the expected enhancement of higher multipole components, reflecting the increased angular resolution of the partial waves.
@@ VII.2 Total scattering cross section
Figure 11: Total scattering cross section in logarithmic scale for a Hernquist black hole at fixed frequency $M\omega=3$. Left panel: Dependence on the density parameter ${\rho_{\text{s}}}M^{2}$, varied from 0 to $0.6$, with the scaled radius fixed at ${r_{\text{s}}}/M=0.5$. Right panel: Dependence on the scaled radius ${r_{\text{s}}}/M$, varied from 0 to $0.6$, with the density parameter fixed at ${\rho_{\text{s}}}M^{2}=0.5$.
The angular dependence of the total scattering cross section, displayed in logarithmic scale as a function of the scattering angle $\theta$, is shown in Fig. 11.
In the left panel, where the scaled radius is fixed at ${r_{\text{s}}}/M=0.5$, increasing the central density parameter ${\rho_{\text{s}}}M^{2}$ leads to a moderate enhancement of the forward scattering peak accompanied by a slight narrowing of the fringe pattern. In contrast, the right panel demonstrates that variations in the scaled radius ${r_{\text{s}}}/M$, for a fixed density ${\rho_{\text{s}}}M^{2}=0.5$, produce a more pronounced effect. In this case, the total scattering cross section increases significantly, and the interference fringes become markedly narrower as ${r_{\text{s}}}/M$ increases from 0 to $0.6$.

### v2

@@ III.1.1 Thermal radiative emission
The contribution to the emitted radiation associated with frequencies around $\omega$ is obtained by restricting attention to the spectral flux within the differential window $\omega,\omega+\mathrm{d}\omega] [$51]. Carrying out this evaluation yields
@@ V Partial wave analysis
Taking into account Eq. (83), the effective potential for the scalar field exhibits a localized barrier that vanishes asymptotically at the event horizon and at spatial infinity. The asymptotic form of the radial wavefunction $\psi(r^{*})$ is therefore constrained: it must be purely ingoing at the horizon, while at infinity it consists of a superposition of incident and reflected waves. Imposing these physical boundary conditions yields  [macedo2015scattering, 79, 80]
@@ VI.1 Low–Frequency Regime
At sufficiently low frequencies, the absorption cross–section of a black hole is known to converge to the area of its event horizon. This universal behavior has been extensively established in the literature and represents a robust feature of black hole scattering processes [85, 86, 87]. In the long–wavelength regime, where the incident wave probes length scales comparable to or larger than the black hole size, the detailed structure of the spacetime becomes subdominant. As a result, the absorption process is governed primarily by the geometric properties of the horizon, causing the cross–section to asymptotically approach the horizon area.
In particular, for scalar field perturbations propagating on a Schwarzschild background, the low–frequency absorption cross–section is analytically shown to coincide with the geometrical area of the event horizon. Remarkably, this limit persists across a wide range of scenarios, including rotating black holes, modified black hole parameters, and nontrivial surrounding geometries [88, 89, 81], provided the incident frequency remains sufficiently small.
Consequently, in the low–frequency regime, the absorption cross–section can be expressed as
@@
The low–frequency behavior of the absorption cross–section for different values of the Hernquist parameters is illustrated in Fig. 7. As expected, when both $r_{\text{s}}$} and $\rho_{\text{s}}$} vanish, the spacetime reduces to the Schwarzschild black hole, for which the low–frequency absorption cross–section approaches $\sigma_{\text{low}}\simeq 16\pi M^{2}$. For a fixed value of ${r_{\text{s}}}/M=0.1$, increasing the dimensionless density parameter ${\rho_{\text{s}}}M^{2}$ from 0 to $0.6$ produces only a negligible modification in the low–frequency absorption cross–section. This weak dependence is clearly visible in the left panel of Fig. 6, where $\sigma_{\text{abs}}$ remains close to the Schwarzschild case, $16\pi M^{2}$, at the low–frequency regime for different values of ${\rho_{\text{s}}}M^{2}$.
In contrast, the absorption cross–section exhibits a more pronounced sensitivity to variations in the scale radius $r_{\text{s}}$}. As shown in the right panel of Fig. 7, increasing ${r_{\text{s}}}/M$ leads to a noticeable enhancement of the absorption cross–section. This trend is consistent with the behavior displayed in Fig. 6, where larger values of ${r_{\text{s}}}/M$ correspond to higher absorption cross–sections at low frequencies.
@@ VI.2 High–Frequency Regime
In this section, we investigate the high–frequency limit of the absorption cross–section, commonly referred to as the geometric cross–section, $\sigma_{\text{geo}}$. In this regime, wave effects become negligible and the absorption process is well described by null geodesic motion. The geometric cross–section is defined as
@@
Substituting Eq. (99) into Eq. (98), the critical impact parameter can be obtained, which in turn allows us to compute the high–frequency absorption cross–section. The final expression for $\sigma_{\text{geo}}$ reads
@@
Figure 6 displays the total absorption cross–section obtained from partial waves with $l=0$ to $l=6$ for different values of the Hernquist parameters. The corresponding geometric cross–section, computed from Eq. (101), is shown by dashed horizontal lines. As expected, the total absorption cross–section approaches a constant value in the high–frequency regime, confirming its convergence toward the geometric cross–section $\sigma_{\text{geo}}$.
Moreover, the high–frequency absorption cross–section exhibits a stronger sensitivity to variations in the scale parameter ${r_{\text{s}}}/M$ than to changes in the density parameter ${\rho_{\text{s}}}M^{2}$. As shown in the left panel of Fig. 6, the modification of the geometric cross–section induced by varying ${\rho_{\text{s}}}M^{2}$ is extremely weak and remains practically indistinguishable. In contrast, the right panel demonstrates that $\sigma_{\text{geo}}$ increases with increasing ${r_{\text{s}}}/M$, indicating that the geometric scale of the system plays a dominant role in determining the high–frequency absorption behavior.
@@ VII Scattering cross section
The observable angular distribution of the scattered radiation is characterized by the differential scattering cross–section, which is related to the scattering amplitude by [93]
@@
The total scattering cross–section is obtained by integrating the differential cross–section over the full solid angle. For a monochromatic wave of frequency $\omega$, this leads to
@@
This expression highlights that the total scattering cross–section arises from a coherent superposition of all partial waves, with each contribution weighted by the corresponding phase shift.
VII.1 Partial scattering cross–section
@@
To investigate the influence of the Hernquist black hole parameters on the scattering properties, we analyze the angular dependence of the differential scattering cross–section for selected multipole numbers. In particular, Figs. 89 display the differential cross–section as a function of the scattering angle $\theta$ for several representative values of $\ell$, illustrating how individual partial waves contribute to the overall scattering pattern.
@@
Figure 8: The normalized partial differential scattering cross–sections of the scalar wave for $M\omega=1$, $\ell=1,~2$, ${r_{\text{s}}}/M=0.3$ and various different Hernquist parameter ${\rho_{\text{s}}}M^{2}$.
@@
Figure 8 presents the partial differential scattering cross-sections for scalar waves with $M\omega=1$ and ${r_{\text{s}}}/M=0.3$, for several choices of ${\rho_{\text{s}}}M^{2}$, comparing the cases $\ell=1$ and $\ell=2$, in the left and right panle, respectively. In both multipole numbers, a clear and consistent pattern emerges: increasing ${\rho_{\text{s}}}M^{2}$ leads to a higher overall amplitude of the partial scattering cross-section. While the $\ell=1$ and $\ell=2$ curves display their usual differences in angular sharpness, their dependence on ${\rho_{\text{s}}}M^{2}$ remains uniform, indicating that the effect of this parameter is broadly insensitive to the multipole structure.
@@
Figure 9: The normalized partial differential scattering cross–sections of the scalar wave for $M\omega=1$, $\ell=1,~2$, ${\rho_{\text{s}}}M^{2}=0.1$ and various different Hernquist parameter ${r_{\text{s}}}/M$.
@@
Figure 9 displays the partial differential scattering cross-sections for scalar waves with fixed parameters $M\omega=1$ and ${\rho_{\text{s}}}M^{2}=0.1$, considering the multipole modes $\ell=1$ and $\ell=2$ in the left and right panels, respectively. For both modes, the qualitative behavior is the same: the partial scattering cross-section increases as the Hernquist parameter ${r_{\text{s}}}/M$ becomes larger. Although the $\ell=1$ and $\ell=2$ curves differ in the shape of their angular patterns, they exhibit the same monotonic dependence on ${r_{\text{s}}}/M$.
@@
Figure 10: The normalized partial scattering cross–sections for different multipole numbers when ${\rho_{\text{s}}}M^{2}={r_{\text{s}}}/M=0.1$.
@@
Figure 10 displays the scalar wave differential scattering cross-sections in the Hernquist black hole spacetime for several multipole numbers, evaluated at ${\rho_{\text{s}}}M^{2}={r_{\text{s}}}/M=0.1$. As the angular momentum mode $\ell$ increases, from $\ell=1$ to $4$, the angular structure of the scattering becomes progressively sharper and more oscillatory, with the dominant contribution concentrated near the forward direction where the cross-section reaches its maximum. This behavior is consistent with the expected enhancement of higher multipole components, reflecting the increased angular resolution of the partial waves.
@@ VII.2 Total scattering cross section
Figure 11: Total scattering cross–section in logarithmic scale for a Hernquist black hole at fixed frequency $M\omega=3$. Left panel: Dependence on the density parameter ${\rho_{\text{s}}}M^{2}$, varied from 0 to $0.6$, with the scaled radius fixed at ${r_{\text{s}}}/M=0.5$. Right panel: Dependence on the scaled radius ${r_{\text{s}}}/M$, varied from 0 to $0.6$, with the density parameter fixed at ${\rho_{\text{s}}}M^{2}=0.5$.
The angular dependence of the total scattering cross–section, displayed in logarithmic scale as a function of the scattering angle $\theta$, is shown in Fig. 11.
In the left panel, where the scaled radius is fixed at ${r_{\text{s}}}/M=0.5$, increasing the central density parameter ${\rho_{\text{s}}}M^{2}$ leads to a moderate enhancement of the forward scattering peak accompanied by a slight narrowing of the fringe pattern. In contrast, the right panel demonstrates that variations in the scaled radius ${r_{\text{s}}}/M$, for a fixed density ${\rho_{\text{s}}}M^{2}=0.5$, produce a more pronounced effect. In this case, the total scattering cross–section increases significantly, and the interference fringes become markedly narrower as ${r_{\text{s}}}/M$ increases from 0 to $0.6$.